OpenStudy (crreedjr):

Mathematical induction...give the counterexample to disprove the following...1^3+2^3+3^3+...+n^3=n^4-n^3+1....I am lost, please help!

OpenStudy (loser66):

If n =3, that is you use 3 terms from the LHS and replace n =3 on the RHS \(1^3+2^3+3^3=1+8+27= 36\) while the RHS is \(3^4-3^3+1=81-27+1=55\) Hence \(LHS\neq RHS\)

OpenStudy (crreedjr):

Why use n=3?

OpenStudy (loser66):

Because it is the case, the equation becomes inequality.